OpenStudy (anonymous):

Find the standard form. Center at origin (0,0); passes through (6,-7)

OpenStudy (anonymous):

circle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if the center is \((0,0)\) you know it is going to look like \(x^2+y^2=r^2\) where \(r^2\) is the square of the distance between \((0,0)\) and \((6,-7)\)

OpenStudy (anonymous):

in other words, \(r^2=6^2+7^2\)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw